OpenStudy (anonymous):

How do I evaluate 8^1/4 * 8^7/4??

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf m^3\times m^5=m ^{3+5}=m^8 }\) can you add you exponents similarly, or need more help? \(\LARGE\color{blue}{ \bf 8^{1/4}\times 8^{7/4}=? }\)

hero (hero):

Hint: \(a^{b/c} a^{d/c} = a^{(b + d)/c}\)

OpenStudy (anonymous):

So it's 8^8/4? Which is simplified to 8^2?

OpenStudy (solomonzelman):

yes ! And 8^2 equals ?

OpenStudy (anonymous):

64! :) thankyou!

OpenStudy (solomonzelman):

You welcome !